UARS Weekly Status Report

01 November 2002


                UARS WEEKLY STATUS REPORT -  1 Nov 2002


We are extremely pleased to report that the UARS Team has apparently
won the 2002 William T. Pecora Award.  This award is presented 
annually to recognize outstanding contributions by individuals or 
groups toward the understanding of the Earth by means of remote 
sensing.  The award is sponsored jointly by the Department of the 
Interior and NASA.  The award to the UARS Team is pending the 
signatures of the Secretary of the Interior and the NASA 
Administrator to be official.
Several of you provided endorsement letters in support of this 
nomination.  Thanks so much for your efforts!
The 2002 William T. Pecora Award will be presented at the 
"Pecora 15 / Land Satellite Information IV Conference" at the 
Adam's Mark Hotel in Denver, CO, during the keynote session 
from 1:00-3:00 p.m. on Tuesday, November 12, 2002.  
Carl (Skip) A. Reber, the original UARS Project Scientist, 
will be accepting the award on the UARS Team's behalf.


Nothing received.


SUSIM resumed operations after the recent anomaly on Oct. 29.
Since then, SUSIM successfully performed its daily mid-resolution
and MgII solar scans during periods of predicted UARS telemetry.


Nothing received.


     The PEM web site has moved.  It has a new URL.  Access the PEM web site
using  The PEM data from the GUVI conjunction
periods has moved to

     PEM continues to work with the GUVI team on the diagnosis of data from
the first conjunction period.  Analysis of the two conjunction passes showed
that the aurora as detected by GUVI was at higher latitude than UARS.  
GUVI images from other times from the first conjunction period show spatial
regions that both GUVI and PEM sampled when aurora was present.  Separation
is several minutes to 10's of minutes when both instruments saw aurora.  A
presentation will be made at the November American Meteorological Society 
meeting on PEM's view of the first conjunction period.  The title is "April 
18-24, 2002: PEM Observations of Atmospheric Energy Input and Ionization
Rates determined from Precipitating Charged Particles".  PEM has seen no aurora 
during the second conjunction period.

     PEM continues to receive NMC data files from NOAA.  No problems have
been encountered.  NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.

*Beginning of power orbit day (PEM activation): array current >5 amps, array 
     voltage >53.5 volts, and battery current >2.0 amps.  
 End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
     array current <9 amps, and battery currents <0.2 amps.
 Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.


Nothing received.


Nothing received.


Nothing received.


WINDII has solved its commanding problems and is now able to operate
normally again. It was most recently activated on October 29, 2002 and
is expected to continue operations until November 14. The Canadian Space
Agency has also agreed to support WINDII operations for one further


The MPG assisted the FOT during analysis of the UARS
spacecraft anomaly which started on October 22.
The MPG continues to serve as the point-of-contact for UARS
instrument operations to support re-engineering of the UARS
Mission Operations Concept to develop capabilities for the
Science Traceability Mission.

Daily processing of instrument Daily Activity Plans continues
with no problems.


Nothing received.

                           REPORTING PERIOD

This  report covers 19 October 2002 (Orbit 60730, GMT Day 292) through
25 October 2002 (Orbit 60834, GMT Day 298).

                         SPACECRAFT OPERATION

The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and   is  performing  nominally  under  two-battery  operations.   The
instrument operational changes were:


       Instrument    Time          Orbit     Comment
       PEM           294/1314      60768     ZEPS/HEPS TO FULLTIME OPS
       PEM           294/1314      60768     POWER ON - AXIS 2
       PEM           294/1830      60771     POWER OFF/ON AXIS 1 & 2
       HALOE         295/0348      60777     POWER OFF
       ALL           295/0419      60777     POWER OFF

The  beta  angle  ranged between the angles of 35.4  degrees  to  62.7
degrees  for  this  report period.  The beta angle is  now  increasing
towards a maximum angle of 71 degrees on 31 October 2002 (DOY 304).

On-Board  Computer  Safehold mode was entered at 295/0122  UTC  on  21
October 2002 during a real-time contact.  The cause of the problem  is
not known at this time but is under investigation.

The OBC controls the High Gain Antenna (HGA) and without the computer,
HGA/TDRS  operations ceased.  UARS nominally operates with transmitter
A  on continuously which is output through the upper omni antenna  for
contingency  situations.  TDRS operations was shifted to this  link  -
UARS omni antenna to TDRS SSA.  This mode of operations is restrictive
for  UARS  since the telemetry is engineering data (reduced scope  and
sampling)  at  a  1  kbps  rate instead of  32  kbps.   The  OBC  dump
capability is reduced to 1 kbps instead of 32 kbps.

OBC  precision pointing control was also lost and UARS was  placed  in
safehold  mode  using  Earth Sensor Assembly-2  (ESA-2).   UARS  earth
sensor  1  has failed and earth sensor 2 is nominally powered  off  to
conserve its remaining useful life.

The  primary focus of the recovery effort was focused on getting  UARS
out  of  OBC safehold and back to precision pointing mode as  soon  as
possible.  All scheduled passes had to be re-scheduled to OMNI  passes
to  allow  maximum contact with UARS.  Then all 16 memory  banks  were
dumped in sequential order to the ground so that an analysis could  be
done to determine what might have initiated the OBC safehold.  Due  to
the  contacts being OMNI SSA contacts the bank dumps were rather  time

The  next course of action was to re-load the OBC with a known  "good"
copy of the UARS hardware load image from the ground.  Again, with the
contacts being OMNI SSA the going was slow, at best.  The FOT made use
of  the  available  HIGH POWER FORWARD feature of TDRS  to  keep  from
having to reload a bank in the case of a dropped NASCOM block.   After
each  bank was loaded it was dumped and compared with what was  loaded
to ascertain that it was exactly what was sent.

The  OBC  was  brought out of safehold on Wednesday  (Day  297)  after
turning  on  all  the  processors  one  by  one.   No  problems   were
encountered  during  this phase of the recovery.   Once  the  OBC  was
running,  the  HGA  tracking was reestablished, and  32  kbps  Science
telemetry along with 32 kbps OBC dump telemetry was restored.

As  soon as the attitude converges, all instruments will be turned  on
as  is  outlined  in the schedule for this flight direction.   In  the
meantime  the  FOT is pursuing a small problem with getting  the  UARS
clock within tolerance for precision science observations.

Spacecraft  battery  1,  2,  and 3 performance  monitoring  continues.
Battery 1 remains OFF the charge relay.  Battery 2 and Battery  3  are
currently maintaining greater than 23.6 V end-of-night (EON) load  bus
voltage.   Battery  temperatures are stable with a  temperature  delta
between Battery 2 and Battery 3 of 0.95 to 0.96 degrees C.  Additional
battery performance data for each battery on 11 October and 18 October
2002 are provided in Appendix D.

The  clock  error  ranged from -?.? to 10.1 msec  during  this  report
period.  Clock rate adjustments are listed in Appendix C.

All  instrument  and subsystem engineers were advised on  the  initial
excursion  of ALERTS or Out-Of-Limits conditions detected by  the  FOT
for  this report.  A summary is listed below and the details of  ALERT
and Out-Of-Limits occurrences are listed in Appendix B.

    OUT-OF-LIMITS  =    00
    ALERTS         =    00

The  following real-time command sheets were executed for  the  listed
operational element(s) during this report period (see Appendix A).

    FOT       =    31
    PEM       =    02
    HALOE     =    01

UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.

    FOT  =    02

Four (4) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).

                             GROUND SYSTEM OPERATION

Two  (2)  operational support problems occurred this period.   Anomaly
Reports are included in Appendix E.

The  data  loss  calculations are no longer  being  supplied  by  Data
Capture Facility (DCF).

                       OTHER SIGNIFICANT EVENTS

                       UARS SATELLITE OPERATIONS

The FOT has focused all of its energy this week on the recovery from
the OBC halt condition.  The OBC Halt occurred at approximately
288/0123 during a regularly scheduled UARS contact.  The cause of this
anomaly is still under investigation and a full Anomaly Investigation
Report (AIR) will be written.  Earth Sensor 2 was brought online at
the next contact through an SSA OMNI TDRS event.  The earth sensor
immediately took out the attitude error that has occurred since the
OBC halt (roughly 8 degrees on roll and yaw).  With the earth sensor
powered on, UARS was in a stable condition and the recovery efforts

The OBC was dumped, reloaded and dumped again over the course of the
next day and a half.  Once the load was confirmed by the dumps, the
OBC was restarted on the 296/1409 event.  All processors were brought
online by the end of the 296/1822 event.  The HGA was put back into
tracking mode over the course of the next few passes and UARS had its
first HGA pass since the anomaly at 296/2209.

The next day was spent watching the UFLTR converge to an attitude
solution using FSS and TAM as inputs.  After observing this
convergence the solution seemed to show higher residuals than those
previous to the anomaly.  After further investigation it was found
that the flight software clock was not properly set.  The FOT uses a
Raytheon written procedure to perform its clock correlation and this
procedure was found to have an error.  All day Friday (DOY 298) was
spent trying to resolve the issue with clock correlation and no
solution was found.  The investigation will continue next week.  Once
this issue is resolved the UARS instruments will be able to be turned
on again.


-     The FOT attended several meetings throughout the week to discuss
  the safehold recovery activities.

                       FUTURE SIGNIFICANT EVENTS

The  next  UARS Yaw slew is scheduled for 18 November 2002 (DOY  322).
This  maneuver will be a Reverse to Forward Flight maneuver  (Southern
Hemisphere viewing).

Attachments:   Appendix A,     Real-time Command Sheets
               Appendix B,     Out-Of-Limits   &   ALERT Occurrences
               Appendix C,     Clock &  Solar  Array  Rate Corrections
               Appendix D,     UARS Battery Performance Data
               Appendix E,     UARS Anomaly Reports

                              APPENDIX A
                       REAL-TIME COMMAND SHEETS
                      Week ending 25 October 2002

ORBIT   AOS        NO.      SUBSYS      REASON
-----   ---        ---      ------      ------
60760   294/0036   02-364   OBC/CLK     CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
60768   294/1314   02-365   PEM         ZEPS/HEPS TO FULLTIME OPS
60768   294/1314   02-366   PEM         POWER ON - AXIS 2
60771   294/1830   02-366   PEM         POWER OFF/ON AXIS 1&2
60776   295/0256   02-367   MACS        EARTH SENSOR 2 ON
60777   295/0348   02-368   OBC         HARDWARE BANK DUMPS
60777   295/0348   02-369   HALOE       POWER OFF
60777   295/0419   02-370   CADH        CDH RECONFIG FOR 1K DUMP THRU OMNI
60777   295/0419   02-371   INST        MANUALLY TURN INST.OFF
60783   295/1400   02-372   OBC         LOAD BANK 0
60783   295/1458   02-373   OBC         LOAD BANK 1
60784   295/1547   02-374   OBC         LOAD BANK 15
60785   295/1631   02-375   OBC         LOAD BANK 14
60785   295/1730   02-376   OBC         BANK 14 DUMP
60790   296/0122   02-383   OBC         LOAD BANK 6
60791   296/0200   02-384   OBC         DUMP BANK 6; LOAD & DUMP BANK7
60791   296/0301   02-385   OBC         LOAD BANK 8
60792   296/0347   02-386   OBC         DUMP BANK 8, LOAD & DUMP BANK9
60792   296/0441   02-387   OBC         LOAD BANK 10
60793   296/0529   02-388   OBC         DUMP BANK 10, LOAD & DUMP BANK11
60794   296/0716   02-389   OBC         LOAD & DUMP BANK 12
60794   296/0803   02-390   OBC         LOAD & DUMP BANK 13
60795   296/0915   02-391   OBC         LOAD & DUMP BANK 14
60795   296/0953   02-392   OBC         LOAD BANK 15
60805   297/0107   02-399   OBC         DISA. MUB FOR HALOE,SUSIM,WINDII
60813   297/1418   02-400   OBC         TABLE 50 LOAD
60814   297/1455   02-401   OBC         MODE 3
60814   297/1601   02-402   OBC         SET GYRO BIAS
60815   297/1650   02-403   OBC         ENA FSS/TAM
60815   297/1742   02-404   OBC         TAM ALIGN
60816   297/1826   02-404   OBC         GYRO BIAS
60816   297/1925   02-404   OBC         GYRO BIAS
60817   297/2007   02-404   OBC         GYRO BIAS
60820   298/0107   02-405   OBC         GYRO BIAS RESET
60821   298/0255   02-405   OBC         GYRO BIAS RESET
60827   298/1229   02-406   OBC         ICAL TO 3 (OPS)
60828   298/1417   02-407   OBC         CLOCK JUMP CMDS
60829   298/1454   02-408   OBC         FSW CLOCK CMDS
60829   298/1559   02-407   OBC         CLOCK JUMP CMDS
60831   298/1923   02-405   OBC         GYRO BIAS RESET
60832   298/2005   02-409   OBC         CLOCK JUMP CMDS

                              APPENDIX B
                      Week ending 25 October 2002

The following Out-Of-Limits occurrences were detected:

--------       -----   -----------                ------
-None this report period.

The following ALERT occurrences were detected:

MNEMONIC     DESCRIPTION                          ORBITS
--------     -----------                          ------
-None this report period.

                              APPENDIX C
                      Week ending 25 October 2002


ORBIT          TIME                   CLOCK RATE CHANGE
-----          -----                  -----------------
60760          294/0059               CLK ADJ FR 0.0 TO 20.6 MSEC/DAY

                              APPENDIX D
                      Week ending 25 October 2002

     18 October 2002 - GMT Day 291 Beta = 34.9 deg, SA Pos = 269 deg

     DIFF V (mV)  TEMP (C)     CURR (amp)  EON V    AVE C/D  AVE
BAT  MAX/MIN      MAX/MIN      MAX/MIN     MIN      RATIO    DOD(%)
---  -------      --------     ----------  -----    -------  ------
1  -151.2/-173.6 -3.68/-4.63  +00.0/+00.0  08.3     0.000     0.0
2   +11.2/-72.8  +4.82/+3.87  +27.2/-14.4  25.6     1.020    19.8
3   +22.4/-22.4  +3.87/+2.91  +26.8/-14.4  25.4     1.020    19.2

     25 October 2002 - GMT Day 298 Beta = 62.7 deg, SA Pos = 269 deg

     DIFF V (mV)  TEMP (C)     CURR (amp)  EON V    AVE C/D  AVE
BAT  MAX/MIN      MAX/MIN      MAX/MIN     MIN      RATIO    DOD(%)
---  -------      --------     ----------  -----    -------  ------
1  -145.6/-168.0 -2.42/-3.68  +00.0/+00.0  08.3     0.000     0.0
2   +00.0/-33.6  +3.55/+2.60  +20.8/-14.4  27.0     1.050    14.9
3   +22.4/-28.0  +2.60/+1.65  +20.0/-14.4  27.0     1.040    14.7

ORB#           AOS TIME   LVL FR-TO
-----          --------   ---------
- None this report period

                              APPENDIX E
                         UARS ANOMALY REPORTS
                      Week ending 25 October 2002

------    ------    -----     ----      ---       -----------------
02-056    FOT       60775     295/0122            OBC CRASH
02-053    FOT       60775     297/1309  30832     GENERIC LATE ACQUISITION

------    ------    -----     ----      ---       -----------------
-None this report period.

------    ------    -----     ----      ---       -----------------
-None this report period.

                         OPEN ANOMALY REPORTS
------   ------    -----   ----      ---     -----------------
02-052   SOLSTICE  60366   267/1951  N/A     LOADING PROBLEM
02-053   FOT       60630   285/1008  30382   GENERIC LATE ACQUISITION
02-054   FOT       60689   289/0819  31607   FORWARD FAIL-OVER AT WHITESAND
                                             20 SEC RTD lost, non-recover
02-056   FOT       60775   295/0122          OBC CRASH